# Conceptualization of this study, Methodology, Software

### v1

@@ 3.2.2 Audio Spectrogram Transformer (AST) architecture
The AST employed in this work follows the base configuration introduced by the pioneer AST model gong2021ast. It adapts the vision transformer (ViT) framework to log-Mel spectrogram inputs, enabling long-range temporal-frequency modeling through multi-head self-attention. Each audio waveform is first converted into a 128-bin log-Mel spectrogram at a $16$ kHz sampling rate. The spectrogram is then partitioned into a non-overlapping $16$ x $16$ time-frequency patches, each one linearly projected into a 768-dimensional embedding space via a convolutional patch-embedding layer. A learnable classification token (CLS) is attached to the sequence, and two-dimensional sinusoidal positional embeddings are added to retain spatial information. The resulting sequence passes through 20 transformer encoder blocks, each composed of a multi-head self-attention module with 20 heads and a feed-forward network of 3072 hidden units with GELU activations. Residual connections, layer normalization, and dropout (rate = 0.1) are applied throughout to enhance generalization. For classification, the final hidden state corresponding to the CLS token is normalized and fed into a linear layer projecting to the 50 classes of the ESC-50 dataset. In the case of fine-tuning, the network is initialized from an Audioset pretrained checkpoint, enabling transfer of acoustic representations of the Audioset dataset. Optimization uses the AdamW optimizer with weight decay of $10^{-4}$, a cosine learning-rate schedule starting at $5.10^{-5}$, and a batch size of $8$. Training proceeds for 10 epochs.
@@ 3.2.3 Retraining the models using UrbanSound8K dataset
To assess the generalization capability of the pre-trained model, we fine-tune it using the UrbanSound8K dataset, a widely used benchmark for environmental sound classification. This retraining process is designed to adapt the ESC-50 pre-trained model to a new dataset while preserving its previously learned feature representations, thus improving its performance on different types of audio data. Instead of training the model from scratch, we utilize transfer learning. This involves leveraging the knowledge embedded in the ESC-50 model and modifying its classification head to align with the specific label structure of UrbanSound8K. By doing so, we efficiently adapt the model to a new domain while retaining valuable learned features from the original training process. The input structure of the model remains unchanged, with retraining applied exclusively to the final classification layer. To achieve optimal adaptation, the model is trained for 50 epochs with a batch size of 32, striking a balance between stability and convergence speed. This training strategy enables the model to refine its decision boundaries for the new dataset while preventing overfitting and ensuring that previously learned knowledge is effectively transferred.
@@ 3.3 Experimental Setup
To evaluate the performance of our proposed approach for environmental sound classification, we conducted a series of experiments using two CNN architectures (CNN-1 and CNN-2) on the UrbanSound8K and ESC-50 datasets. The UrbanSound8K dataset contains 8,732 labeled audio clips (up to 4 seconds each) spanning 10 environmental sound classes (e.g., car horn, siren, dog bark), and follows a 10-fold cross-validation structure. The ESC-50 dataset consists of 2,000 five-second clips organized into 50 balanced classes. Our experimental pipeline involves four main stages: data preprocessing, feature extraction, model training (pre-training and fine-tuning), and performance evaluation. All audio clips were resampled to 22,050 Hz and trimmed or zero-padded to 3 seconds. Feature values were normalized to zero mean and unit variance. Each clip was transformed into multiple time-frequency representations, including MFCC, GTCC, Chroma, Tonnetz, Spectral Contrast, and Log-Mel Spectrogram. These features were resized to a uniform dimension of 128 × 128 using zero-padding or interpolation, and then stacked to form multi-channel inputs analogous to RGB images. For training, each CNN model was initially pre-trained on ESC-50 and fine-tuned on UrbanSound8K using transfer learning. During fine-tuning, pre-trained convolutional layers were frozen and only the final classification layers were retrained. The models were trained using categorical cross-entropy loss and the Adam optimizer, with accuracy as the main evaluation metric. Training was conducted for 150 epochs on ESC-50 and 50 epochs on UrbanSound8K, using a batch size of 32. Validation was performed using held-out test sets to evaluate generalization. All experiments were carried out on Google Colab, leveraging its NVIDIA H100 GPU for accelerated computation. The models were implemented using the Keras API with TensorFlow as the backend, and Librosa was employed for audio signal processing and feature extraction.
@@ 4 Results and Analysis
Figure 1 illustrates the training and validation accuracy and loss of the CNN-1 and CNN-2 models during fine-tuning on the UrbanSound8K dataset, following pre-training on ESC-50, when MFCC + GTCC + CH + LM features are used.
@@ 4.1 Performance Trends and Observations
In this subsection, we provide a comprehensive analysis of the experimental findings, focusing on the performance behavior of the proposed models under various conditions. As shown in Table 1, the column "Training Setup" indicates the transfer learning strategy used during training. “All.L” refers to training all convolutional layers from scratch (i.e., no layers are frozen), and “Last.L” refers to fine-tuning only the last layer of the network while freezing the rest, which allows for transfer learning from pretrained weights on the ESC-50 dataset to the UrbanSound8K dataset.
@@ 4.2 Comparison with AST Model
The comparative evaluation between the CNN and the AST models reveals a clear performance distinction across different training configurations, as demonstrated in Table 2. When trained on the ESC-50 dataset from scratch (All.L), the AST model using a single Mel-Spectrogram input achieved a validation accuracy of 43%, lower than the results obtained by both CNN models 1 and 2, which were 68% (using LM) and 66% (using LM+TZ), respectively. The performance gap underscores the AST model’s sensitivity to limited training data, as its self-attention mechanism typically requires larger datasets or extensive pretraining to reach higher accuracy gong2021ast.
@@
The performance difference can be attributed to both architectural and data factors. CNNs, particularly when supplied with stacked features, effectively capture complementary time-frequency representations and additional representations when aggregated with other features, enabling them to generalize well even with moderate training data sizes. The AST, while capable of modeling long-range dependencies, is more reliant on a large-scale pretraining dataset to fully exploit its attention mechanisms. We evaluated the AST with Mel-Spectrogram input using AudioSet pretraining followed by ESC fine-tuning, which achieved a validation accuracy of 99%, confirming the boost that large-scale pretraining can provide. However, without large-scale pretraining, the AST struggles to match the capabilities of CNNs.
@@ 4.3 Training Time and Computational Efficiency Analysis
The computational efficiency of deep learning models is a key factor in real-world applications, particularly for deployment on resource-constrained or real-time systems. In our experiments, we analyzed the training time and inference efficiency of the CNN-1 and CNN-2 models fine-tuned using a last-layer retraining strategy. The comparison in terms of model size and inference time per sample is summarized in Table3.
As shown in Table 3, CNN-1 has fewer parameters (146K) compared to CNN-2 (151K), reflecting its slightly simpler architecture. In terms of inference efficiency, CNN-1 achieved a mean inference time of 18.13 ms per sample with a standard deviation of 2.21 ms, while CNN-2 exhibited a higher inference cost, with a mean inference time of 29.82 ms and a standard deviation of 2.83 ms. This difference is primarily attributed to the inclusion of Batch Normalization layers in CNN-2, which introduce additional computations during inference despite their benefits for training stability. Regarding training efficiency, CNN-1 consistently converged faster than CNN-2, requiring fewer training iterations to reach its best validation performance. The simpler architecture of CNN-1 enables more efficient gradient updates and lower computational overhead during training. In contrast, CNN-2 demonstrated improved training stability and smoother convergence behavior, benefiting from the regularization effect of Batch Normalization, particularly in later training stages.

### v2

@@ 3.3.2 Audio Spectrogram Transformer (AST) architecture
The AST employed in this work follows the base configuration introduced by the pioneer AST model [9]. It adapts the vision transformer (ViT) framework to log-Mel spectrogram inputs, enabling long-range temporal-frequency modeling through multi-head self-attention. Each audio waveform is first converted into a 128-bin log-Mel spectrogram at a 16 kHz sampling rate. The spectrogram is then partitioned into a non-overlapping 16 x 16 time-frequency patches, each one linearly projected into a 768-dimensional embedding space via a convolutional patch-embedding layer. A learnable classification token (CLS) is attached to the sequence, and two-dimensional sinusoidal positional embeddings are added to retain spatial information. The resulting sequence passes through 20 transformer encoder blocks, each composed of a multi-head self-attention module with 20 heads and a feed-forward network of 3072 hidden units with GELU activations. Residual connections, layer normalization, and dropout (rate = 0.1) are applied throughout to enhance generalization.
For classification, the final hidden state corresponding to the CLS token is normalized and fed into a linear layer projecting to the 50 classes of the ESC-50 dataset. In the case of fine-tuning, the network is initialized from an Audioset pretrained checkpoint, enabling transfer of acoustic representations of the Audioset dataset. Optimization uses the AdamW optimizer with weight decay of $10^{-4}$, a cosine learning-rate schedule starting at $5\times 10^{-5}$, and a batch size of 8. Training proceeds for 10 epochs.
@@ 3.3.3 Retraining the models using UrbanSound8K dataset
To assess the generalization capability of the pre-trained model, we fine-tune it using the UrbanSound8K dataset, a widely used benchmark for environmental sound classification. This retraining process is designed to adapt the ESC-50 pre-trained model to a new dataset while preserving its previously learned feature representations, thus improving its performance on different types of audio data. Instead of training the model from scratch, we utilize transfer learning. This involves leveraging the knowledge embedded in the ESC-50 model and modifying its classification head to align with the specific label structure of UrbanSound8K. By doing so, we efficiently adapt the model to a new domain while retaining valuable learned features from the original training process.
The input structure of the model remains unchanged, with retraining applied exclusively to the final classification layer. This selective fine-tuning approach ensures that the lower-level feature extraction layers, which have already learned meaningful audio representations, remain intact, allowing the model to generalize well across different datasets. To achieve optimal adaptation, the model is trained for 50 epochs with a batch size of 32, striking a balance between stability and convergence speed. This training strategy enables the model to refine its decision boundaries for the new dataset while preventing overfitting and ensuring that previously learned knowledge is effectively transferred.
@@ 3.4 Experimental Setup
To evaluate the performance of our proposed approach for environmental sound classification, we conducted a series of experiments using two CNN architectures (CNN-1 and CNN-2) on the UrbanSound8K and ESC-50 datasets. The UrbanSound8K dataset contains 8,732 labeled audio clips (up to 4 seconds each) spanning 10 environmental sound classes (e.g., car horn, siren, dog bark), and follows a 10-fold cross-validation structure. The ESC-50 dataset consists of 2,000 five-second clips organized into 50 balanced classes.
Our experimental pipeline involves four main stages: data preprocessing, feature extraction, model training (pre-training and fine-tuning), and performance evaluation. All audio clips were resampled to 22,050 Hz and trimmed or zeropadded to 3 seconds. Feature values were normalized to zero mean and unit variance. Each clip was transformed into multiple time-frequency representations, including MFCC, GTCC, Chroma, Tonnetz, Spectral Contrast, and Log-Mel Spectrogram. These features were resized to a uniform dimension of 128 × 128 using zero-padding or interpolation, and then stacked to form multi-channel inputs analogous to RGB images. For training, each CNN model was initially pre-trained on ESC-50 and fine-tuned on UrbanSound8K using transfer learning. During fine-tuning, pre-trained convolutional layers were frozen and only the final classification layers were retrained. The models were trained using categorical cross-entropy loss and the Adam optimizer, with accuracy as the main evaluation metric. Training was conducted for 150 epochs on ESC-50 and 50 epochs on UrbanSound8K, using a batch size of 32. Validation was performed using held-out test sets to evaluate generalization.
All experiments were carried out on Google Colab, leveraging its NVIDIA H100 GPU for accelerated computation. The models were implemented using the Keras API with TensorFlow as the backend, and Librosa was employed for audio signal processing and feature extraction. To further assess the effect of different feature combinations, we conducted 5-fold cross-validation experiments using both CNN architectures on the ESC-50 dataset. The results, presented in Table 2, report classification accuracy across each fold along with the average performance.
Table 2: Accuracy of CNN-1 and CNN-2 across 5 folds with different feature combinations.
Models Features Fold 1 Fold 2 Fold 3 Fold 4 Fold 5 Average Accuracy
CNN-1 LM 0.70 0.67 0.69 0.66 0.66 0.68
LM+TZ 0.66 0.66 0.68 0.62 0.63 0.65
LM+MFCC 0.65 0.64 0.66 0.64 0.61 0.64
MFCC+TZ 0.65 0.59 0.61 0.60 0.64 0.62
LM+SPC+CH 0.61 0.59 0.63 0.62 0.65 0.62
MFCC+GTCC+CH+LM 0.69 0.67 0.65 0.66 0.68 0.67
CNN-2 LM 0.43 0.47 0.45 0.46 0.44 0.45
LM+TZ 0.65 0.62 0.69 0.68 0.66 0.66
LM+MFCC 0.61 0.62 0.57 0.60 0.58 0.59
MFCC+TZ 0.62 0.59 0.63 0.62 0.65 0.62
LM+SPC+CH 0.52 0.56 0.54 0.51 0.53 0.53
MFCC+GTCC+CH+LM 0.59 0.56 0.59 0.57 0.60 0.58
@@ 4 Results and Analysis
Figure 4 to Figure 9 illustrates the training and validation accuracy and loss of the CNN-1 and CNN-2 models during fine-tuning on the UrbanSound8K dataset, following pre-training on ESC-50, when different input features are used.
@@
Figure 4: Comparison of Training and Validation Accuracy and Loss for CNN-1 and CNN-2 using Log-Mel (LM) Features Only.
Figure 5: Comparison of Training and Validation Accuracy and Loss for CNN-1 and CNN-2 using stacked Log-Mel (LM) and MFCC features.
Figure 6: Comparison of Training and Validation Accuracy and Loss for CNN-1 and CNN-2 using stacked Log-Mel (LM) and Tonnetz (TZ) features.
Figure 7: Comparison of Training and Validation Accuracy and Loss for CNN-1 and CNN-2 using stacked MFCC and TZ features.
Figure 8: Comparison of Training and Validation Accuracy and Loss for CNN-1 and CNN-2 using stacked Log-Mel (LM), Spectral Contrast (SPC), and Chroma (CH) features.
@@ 4.1 Performance Trends and Observations
In this subsection, we provide a comprehensive analysis of the experimental findings, focusing on the performance behavior of the proposed models under various conditions. As shown in Table 3, the column "Training Setup" indicates the transfer learning strategy used during training. “All.L” refers to training all convolutional layers from scratch (i.e., no layers are frozen) and “Last.L” refers to fine-tuning only the last layer of the network while freezing the rest, which allows for transfer learning from pretrained weights on the ESC-50 dataset to the UrbanSound8K dataset. A significant performance improvement is observed when moving from the ESC, All.L setup to the ESC + US8K, Last.L setup. This improvement is primarily due to the transfer learning strategy and the increased diversity of training data rather than architectural modifications alone. In the ESC, All.L configuration, the networks are trained from scratch using only ESC-50 data, which limits the amount of available acoustic variability. In contrast, the ESC + US8K, Last.L configuration leverages pretrained representations learned from ESC-50 and adapts them to UrbanSound8K through fine-tuning of the final layer. This cross-dataset transfer exposes the models to a broader distribution of environmental sounds, enabling more robust feature representations and better generalization. Therefore, the observed accuracy gains are largely attributable to the expanded data regime and transfer-learning strategy rather than solely to architectural differences.
@@ 4.2 Comparison with AST Model
The comparative evaluation between the CNN and the AST models reveals a clear performance distinction across different training configurations, as demonstrated in Table 4. When trained on the ESC-50 dataset from scratch (All.L), the AST model using a single Mel-Spectrogram input achieved a validation accuracy of 43%, lower than the results obtained by both CNN models 1 and 2, which were 68% (using LM) and 66% (using LM+TZ), respectively. The performance gap underscores the AST model’s sensitivity to limited training data, as its self-attention mechanism typically requires larger datasets or extensive pretraining to reach higher accuracy  [9].
@@
The performance difference can be attributed to both architectural and data factors. CNNs, particularly when supplied with stacked features, effectively capture complementary time-frequency representations and additional representations when aggregated with other features, enabling them to generalize well even with moderate training data sizes. The AST, while capable of modeling long-range dependencies, is more reliant on a large-scale pretraining dataset to fully exploit its attention mechanisms. We evaluated the AST with Mel-Spectrogram input using AudioSet pretraining followed by ESC fine-tuning, which achieved a validation accuracy of 99%, confirming the boost that large-scale pretraining can provide. However, without the large-scale pretraining, AST struggles to match the capabilities of CNNs.
@@ 4.3 Training Time and Computational Efficiency Analysis
The computational efficiency of a deep learning model is a key factor in real-world applications, particularly for deployment on resource-constrained or real-time systems. In our experiments, we analyzed the training time and inference efficiency of the CNN-1 and CNN-2 models when fine-tuned on the UrbanSound8K dataset using a last-layer retraining strategy.
As shown in Table 5, CNN-1 has fewer parameters (146K) compared to CNN-2 (151K), reflecting its slightly simpler architecture. In terms of inference efficiency, CNN-1 achieved a mean inference time of 21.92 ms per sample with a standard deviation of 2.15 ms, while CNN-2 exhibited a higher inference cost, with a mean inference time of 30.95 ms and a standard deviation of 4.07 ms. These results were obtained over 50 inference iterations per sample to ensure statistical reliability. This difference is primarily attributed to the inclusion of Batch Normalization layers in CNN-2, which introduce additional computations during inference despite their benefits for training stability. Regarding training efficiency, CNN-1 consistently converged faster than CNN-2, requiring fewer training iterations to reach its best validation performance. The simpler architecture of CNN-1 enables more efficient gradient updates and lower computational overhead during training. In contrast, CNN-2 demonstrated improved training stability and smoother convergence behavior, benefiting from the regularization effect of Batch Normalization, particularly in later training stages.